In two years Orgo will be twice as old as he was 5 years ago how old is he now?
Answer by richwmiller(17219)

x+2=2*(x-5)
x+2=2x-10
subtract x
2=x-10
add 10
12=x
check
14=2*(7)
14=14
ok